# Shamier Settle, Matthew Shapiro, Emerita Torres, and Anand Vidyashankar for discussion and feedback on the work described

### v2

@@ Abstract
We estimate the number of street vendors in New York City. First, we summarize the process by which vendors receive licenses and permits to operate legally in New York City. Second, we describe a survey that was administered by the Street Vendor Project while distributing Coronavirus relief aid to vendors operating in New York City both with and without a license or permit. Third, we calculate the total number of vendors using ratio estimation, finding approximately 23,000 street vendors operate in New York City—20,500 mobile food vendors and 2,300 general merchandise vendors—with one third located in just six ZIP Codes—11368 (16%), 11372 (3%), and 11354 (3%) in North and West Queens and 10036 (5%), 10019 (4%), and 10001 (3%) in the Chelsea and Clinton neighborhoods of Manhattan. Fourth, we provide a theoretical justification of our estimates based on the theory of point processes. Finally, we evaluate the accuracy of the ratio estimator when the distribution of vendors is explained by a Poisson or Yule process, and we discuss several policy implications. In particular, our estimates suggest the American Community Survey misses the majority of New York City street vendors.
@@ 1. Introduction
Street vendors are New York City’s smallest businesses, selling food and merchandise from carts, stalls, and trucks throughout the five boroughs. They are an iconic part of the urban landscape and a thriving sector of the local economy, contributing millions of dollars in government revenue through taxes, fines, and fees. [^1] Fees from mobile food vending licenses and permits provide nearly a million dollars in revenue to New York City each year (Mosher and Turnquist 2024). Fines paid by vendors provide approximately \$200,000 each year according to data from the New York City Office of Administrative Trials and Hearings. Perhaps most importantly, street vending historically benefits underserved communities, both because vendors operate in neighborhoods with limited access to traditional stores and because vending is one of a handful of occupations in which New Yorkers of all backgrounds, immigrants in particular, are able to achieve economic mobility and a chance at the American dream (Burrows and Wallace 1998, chap. 42).
@@
In this paper, we propose an estimate for the number of street vendors in New York City, including those that operate without a license or permit. We present our work in four sections. In Section 2, we review the process by which vendors receive permits and licenses, and we describe a survey administered by the Street Vendor Project at the Urban Justice Center while distributing Coronavirus relief aid to vendors operating both with and without a license or permit.
In Section 3, we use the fact that the number of licenses and permits are limited by law to construct a ratio estimator for the number of street vendors. We find approximately 23,000 street vendors operate in New York City: 20,500 mobile food vendors and 2,300 general merchandise vendors. One third are located in just six ZIP Codes: 11368 (16%), 11372 (3%), and 11354 (3%) in North and West Queens and 10036 (5%), 10019 (4%), and 10001 (3%) in the Chelsea and Clinton neighborhoods of Manhattan. We then provide a theoretical justification of our estimates, assuming the spatial distribution of survey respondents is representative and well-approximated by a family of inhomogeneous Poisson processes.
In Section 4, we examine the validity of these assumptions. We evaluate whether the spatial distribution of respondents is representative by comparing it to the American Community Survey and administrative data, and we evaluate the Poisson process assumption by considering a more general model in which the arrival of street vendors follows a pure birth process, of which the Poisson and Yule processes are special cases.
@@
2. We consider two types of vendors: food vendors and merchandise vendors
@@
We limit our analysis to two types of street vendors: mobile food vendors—vendors who sell food items such as sandwiches, drinks, and cut fruit—and general merchandise vendors—vendors who sell merchandise items such as electronics, clothing, and accessories. We refer to these vendors as food vendors and merchandise vendors, respectively. In Section 2.1, we review the process by which vendors receive permits and licenses as required by law. In Section 2.2, we describe a survey administered by the Street Vendor Project while distributing Coronavirus relief aid to vendors operating in New York City both with and without a license or permit.
2.1 The number of licenses and permits are limited by local law
@@
The final exception are concessionaires that operate on New York City parkland through an NYC Parks permit or license. According to the New York City Department of Parks and Recreation, there are approximately 400 concessions within New York City parks. Many offer food services, ranging from food carts to restaurants. We do not consider First Amendment vendors or park concessionaires in this paper. Our estimates are limited to food and merchandise vendors operating outside of parks.
2.2 The Street Vendor Project surveyed two thousand street vendors in 2021
@@
The survey included 100 questions and was conducted in eight different languages: Arabic, Bangla, Cantonese, English, French, Mandarin, Spanish, Tibetan, and Wolof. [^4] Vendors who spoke other languages, such as Dari, Farsi, Russian, and Turkish, typically spoke English with a high level of fluency, and those surveys were conducted in English. Similarly, vendors who spoke indigenous languages from Latin America, such as Nahuatl or Quechua, often spoke Spanish with a high level of fluency. The survey items solicited a variety of information from vendors, such as logistical information (e.g., vending location, residential location, and frequency of operation), economic information (e.g., items sold, income, and expenses), and demographic information (e.g., age, race, ethnicity, and immigration status). Most relevant is the fact that respondents classified themselves by the goods they sold (e.g., food vendors, merchandise vendors, First Amendment vendors, etc.), and respondents indicated whether they had the relevant licenses and permits to vend. Of the 2,060 responses, 1,400 identified as food vendors and 559 as merchandise vendors. The remaining 101 respondents were predominantly First Amendment Vendors, which we exclude from our analysis.
@@
Of the 1,400 food vendors, 349 (25%) indicated they had a permit to vend. Of the 559 merchandise vendors, 505 were not veterans, of which 308 (61%) indicated they had a license to vend. The number of respondents is listed by neighborhood in Section 7.1, Table 1. (Neighborhoods with few respondents are grouped together.) A map of the number of respondents by ZIP Code Tabulation Area (ZIP Code) is shown in the top left panel of Section 7.2, Figure 1. Approximately five percent of vendors did not identify their vending location. These vendors are excluded from Table 1 and Figure 1, although doing so does not meaningfully change the estimated number of vendors that operate in New York City. Veteran merchandise vendors are also excluded from Table 1 and Figure 1, although these vendors are reflected in our city-wide estimate of 23,000 vendors.
3. We calculate the number of street vendors using ratio estimation
@@
We estimate approximately 23,000 street vendors operate in New York City: 20,500 mobile food vendors and 2,300 general merchandise vendors. We arrive at the estimate by using ratio estimation, which leverages the fact that the number of licenses and permits are limited by local law.
@@
In Section 3.1, we provide a simple explanation of ratio estimation, which is intended to be accessible to a general readership. In Section 3.2, we provide a theoretical justification of our estimates based on the theory of point processes, which, while more technical, is intended to provide a somewhat general argument for ratio estimation with spatial data when a design-based approach cannot be justified. Inference is conducted using increasing domain asymptotics, and the details are outlined in Section 7.3.
3.1 We estimate 23,000 street vendors operate in New York City
Ratio estimation is a common approach for estimating the size of a population. See Cochran (1978) and Hald (1998, chap. 16) for a historical discussion and Lohr (2021, chap. 4) for an introduction. The popularity of ratio estimation is due in large part to its simplicity: Among the $1,400$ food vendors surveyed, roughly a quarter had a permit. Suppose this ratio is representative in that one quarter of all food vendors have a permit. It follows that the total number of food vendors is $20,500\approx 5,100/(1/4)$ because one quarter of $20,500$ is approximately $5,100$, the number of permits available.
By similar reasoning—i.e., exploiting the fact that the number of general merchandise licenses is limited for non-veterans—the number of non-veteran merchandise vendors is estimated to be approximately $1,300$. When combined with the estimated $1,000$ veteran merchandise vendors (see Section 2.1) and $20,500$ food vendors, the total number of street vendors is approximately $23,000$.
The main assumption of ratio estimation is that the survey data are representative in the sense that, were survey operations to continue, the percentage of respondents with a permit (in the case of food vendors) or a license (in the case of merchandise vendors) would remain the same. If we further assume that the spatial distribution of respondents is representative, we can also estimate the total number of vendors by neighborhood. [^5] We use the same ratio for each neighborhood. Ratio estimation by neighborhood is not possible since only the locations of the vendors who participated in the survey are known. These estimates are listed in Section 7.1, Table 1. We find that approximately one quarter of street vendors are located in the Chelsea and Clinton neighborhoods of Manhattan. Another quarter is located in West Queens and Lower Manhattan.
We check our estimates by comparing them to a second, independent assessment conducted by the Street Vendor Project in the Bronx on May 13 and 15, 2022, in which the number of visible street vendors was documented through on-the-ground observations. The independent assessment identified 188 total street vendors in the Fordham area (Fordham Road BID and Street Vendor Project 2024), whereas Table 1 indicates 206 vendors in Bronx Park and Fordham—an estimate approximately ten percent higher. To demonstrate that this difference can be explained by sampling variation, we derive the standard error of the ratio estimate in Section 3.2.
3.2 We model the location of street vendors as a Poisson process
@@
We provide a point process justification of the ratio estimation procedure described in Section 3.1. The purpose is to examine the underlying assumptions and derive the standard error of the estimate.
@@
Let $\{\Pi_{i}\}$ denote a family of inhomogeneous spatial Poisson processes, referencing the location $x\in A\subset\mathbb{R}^{2}$ of each street vendor at the time the survey was conducted. The index $i$ denotes the status of the vendor. To fix ideas, let $i=0$ indicate mobile food vendors with a permit, and let $i=1$ indicate mobile food vendors without a permit.
@@
The Poisson process assumption may be justified by the law of rare events. New York City can be partitioned into a large number of theoretically vendable locations. Whether a vendor is located within a partition has a vanishingly small probability, such that the number of vendors in any area is well approximated by a Poisson distribution. This approximation is accurate even if these probabilities are weakly dependent. For example, see Freedman (1974), Chen (1975), and Serfling (1975). Note that other justifications are possible. For example, see Section 4.2 for a justification of the Poisson process based on the theory of birth processes.
@@
so that by conditioning on $N_{1}(A)$, we arrive at the following probability model for the number of respondents with and without permits
$n_{0}(A)\sim$ $\ \text{Poisson}\big{(}p\,\theta\big{)}$
$n_{1}(A)\ |\ N_{1}(A)\sim$ $\ \text{Binomial}\big{(}p,\,N_{1}(A)\big{)}$
The maximum likelihood estimates for $p$ and $\theta$ are
$\hat{p}=\dfrac{n_{1}(A)}{N_{1}(A)}\qquad\text{and}\qquad\hat{\theta}=\dfrac{N_{1}(A)\,n_{0}(A)}{n_{1}(A)}$
@@
The estimator $\hat{\theta}$ is called the ratio estimator because it depends on the ratio of random variables $n_{0}(A)\,/\,n_{1}(A)$. (The total number of permitted vendors, $N_{1}(A)$, is considered fixed.) Using increasing domain asymptotics, the estimator is asymptotically normal with mean $\theta$ and standard error
$\text{SE}(\hat{\theta})=\theta\sqrt{\dfrac{1}{p\,\theta}+\dfrac{1-p}{p\,N_{1}(A)}}$
@@
Substituting $\hat{p}$ and $\hat{\theta}$ for $p$ and $\theta$ yields the following plug-in estimator for the standard error,
$\hat{\text{SE}}(\hat{\theta})=\dfrac{N_{1}(A)\,n_{0}(A)}{n_{1}(A)}\sqrt{\dfrac{1}{n_{0}(A)}+\dfrac{1}{n_{1}(A)}-\dfrac{1}{N_{1}(A)}}$
See Section 7.3 for details. The margin of error reported in Section 7.1, Table 1 is two standard errors, half of the width of a 95% confidence interval. Note that the margin of error for the 206 vendors estimated in Bronx Park and Fordham is 44, and thus this estimate is consistent with the 188 street vendors counted independently by Fordham Road BID and Street Vendor Project (2024) as described in Section 3.1.
4. We examine the validity of our assumptions using additional information

### v4

@@ Abstract
We estimate the number of street vendors in New York City. First, we summarize the process by which vendors receive licenses and permits to operate legally in New York City. Second, we describe a survey that was administered by the Street Vendor Project while distributing Coronavirus relief aid to vendors operating in New York City both with and without a license or permit. Third, we review ratio estimation and provide a theoretical justification based on the theory of point processes. Fourth, we use ratio estimation to calculate the total number of vendors, finding approximately 23,000 street vendors operate in New York City—20,500 mobile food vendors and 2,400 general merchandise vendors—with one third located in just six ZIP Codes—11368 (16%), 11372 (3%), and 11354 (3%) in North and West Queens and 10036 (5%), 10019 (4%), and 10001 (3%) in the Chelsea and Clinton neighborhoods of Manhattan. Finally, we evaluate the accuracy of the ratio estimator when the distribution of vendors is explained by a Poisson or Yule process, and we discuss several policy implications. In particular, our estimates suggest the American Community Survey misses the majority of New York City street vendors.
@@ 1. Introduction
Street vendors are New York City’s smallest businesses, selling food and merchandise from carts, stalls, and trucks throughout the five boroughs. They are an iconic part of the urban landscape and a thriving sector of the local economy, contributing millions of dollars in government revenue through taxes, fines, and fees. [^1] Fees from mobile food vending licenses and permits provide nearly a million dollars in revenue to New York City each year (Mosher and Turnquist 2024). Fines paid by vendors provide approximately \$200,000 each year according to data from the New York City Office of Administrative Trials and Hearings. Perhaps most importantly, street vending historically benefits underserved communities, both because vendors operate in neighborhoods with limited access to traditional stores and because vending is one of a handful of occupations in which New Yorkers of all backgrounds, immigrants in particular, can achieve economic mobility and a chance at the American dream (Burrows and Wallace 1998, chap. 42).
@@
In this paper, we propose an estimate for the number of street vendors in New York City, including those that operate without a license or permit. We present our work in four sections. In Section 2, we review the process by which vendors receive licenses and permits, and we describe a survey administered by the Street Vendor Project at the Urban Justice Center while distributing Coronavirus relief aid to vendors operating both with and without a license or permit.
In Section 3, we use the fact that the number of licenses and permits are limited by law to construct a ratio estimator for the number of street vendors. We then provide a theoretical justification of our estimates, assuming the spatial distribution of survey respondents is representative and well-approximated by a set of inhomogeneous Poisson processes. We find approximately 23,000 street vendors operate in New York City: 20,500 mobile food vendors and 2,400 general merchandise vendors. One third are located in just six ZIP Codes: 11368 (16%), 11372 (3%), and 11354 (3%) in North and West Queens and 10036 (5%), 10019 (4%), and 10001 (3%) in the Chelsea and Clinton neighborhoods of Manhattan. We also derive standard errors.
In Section 4, we examine the validity of our assumptions. We evaluate whether the spatial distribution of respondents is representative by comparing it to the American Community Survey and administrative data, and we evaluate the Poisson process assumption by considering a more general model in which the arrival of street vendors follows a pure birth process, of which the Poisson and Yule processes are special cases. The Yule process may be appropriate if street vendors cluster within markets, in which case the data might have extra-Poisson variation. We rederive the standard errors from Section 3 to account for this variation.
@@
2. Background and Survey Design
@@
We limit our analysis to two types of street vendors: mobile food vendors—vendors who sell food items such as sandwiches, drinks, and cut fruit—and general merchandise vendors—vendors who sell merchandise items such as electronics, clothing, and accessories. We refer to these vendors as food vendors and merchandise vendors, respectively. In Section 2.1, we review the process by which vendors receive licenses and permits as required by law. In Section 2.2, we describe a survey administered by the Street Vendor Project while distributing Coronavirus relief aid to vendors operating in New York City both with and without a license or permit.
2.1 Background
@@
The final exception are concessionaires that operate on New York City parkland through an NYC Parks license or permit. According to the New York City Department of Parks and Recreation, there are approximately 400 concessions within New York City parks. Many offer food services, ranging from food carts to restaurants. We do not consider First Amendment vendors or park concessionaires in this paper. Our estimates are limited to food and merchandise vendors operating outside of parks.
2.2 Survey Design
@@
The survey included 100 questions and was conducted in nine different languages: Arabic, Bangla, Cantonese, English, French, Mandarin, Spanish, Tibetan, and Wolof. [^4] Vendors who spoke other languages, such as Dari, Farsi, Russian, and Turkish, typically spoke English with a high level of fluency, and those surveys were conducted in English. Similarly, vendors who spoke indigenous languages from Latin America, such as Nahuatl and Quechua, typically spoke Spanish with a high level of fluency. The survey items solicited a variety of information from vendors, such as logistical information (e.g., vending location, residential location, and frequency of operation), economic information (e.g., items sold, income, and expenses), and demographic information (e.g., age, race, ethnicity, and immigration status). Most relevant is the fact that respondents classified themselves by the goods they sold (e.g., food vendors, merchandise vendors, First Amendment vendors, etc.), and respondents indicated whether they had the relevant licenses and permits to vend. Of the 2,060 responses, 1,400 identified as food vendors and 559 as merchandise vendors. The remaining 101 respondents were predominantly First Amendment Vendors, which we exclude from our analysis.
@@
Of the 1,400 food vendors, 349 (25%) indicated they had a permit to vend. Of the 559 merchandise vendors, 505 were not veterans, of which 308 (61%) indicated they had a license to vend. The number of respondents is listed by neighborhood in Section 7.1, Table 1. (Neighborhoods with few respondents are grouped together.) A map of the number of respondents by ZIP Code Tabulation Area (ZIP Code) is shown in the top left panel of Section 7.2, Figure 1. Approximately five percent of vendors did not identify their vending location. These vendors are included in the New York City total, but they are excluded from the neighborhood estimates in Table 1 and Figure 1. (For this reason, the Respondent and Population columns do not sum exactly to the New York City total.) Veteran merchandise vendors are excluded entirely from Table 1 and Figure 1, although these vendors are reflected in our overall estimate of 23,000 vendors.
3. Methodology and Data Analysis
@@
We estimate approximately 23,000 street vendors operate in New York City: 20,500 mobile food vendors and 2,400 general merchandise vendors. We arrive at the estimate by using ratio estimation, which leverages the fact that the number of licenses and permits are limited by local law.
@@
In Section 3.1, we provide a simple explanation of ratio estimation, which is intended to be accessible to a general readership. In Section 3.2, we provide a theoretical justification of our estimates based on the theory of point processes, which, while more technical, is intended to provide a somewhat general argument for ratio estimation with spatial data when a design-based approach cannot be justified. Inference is conducted using increasing domain asymptotics, and the details are outlined in Section 7.3. In Section 3.3, we apply the methodology to the survey data discussed in Section 2.2.
3.1 Ratio Estimation
Ratio estimation is a common approach for estimating the size of a population. See Cochran (1978) and Hald (1998, chap. 16) for a historical discussion and Lohr (2021, chap. 4) for an introduction. We provide a simple explanation based on cross-multiplication, also called the rule of three. The purpose is to highlight the main assumption. A more rigorous derivation is provided in Section 3.2.
Consider a fixed region $A$, and let $N_{i}(A)$ denote the number of individuals in region $A$ of type $i$. To fix ideas, let $A$ be New York City, let $i=0$ indicate mobile food vendors with a permit, and let $i=1$ indicate mobile food vendors without a permit. Also, let $n_{i}(A)$ denote the number of respondents of type $i$.
We observe $N_{1}(A)$, $n_{1}(A)$, and $n_{0}(A)$ from which we estimate $\theta(A)=\mathbb{E}\left[N_{0}(A)\right]$, the expected number of vendors without a permit. The expected number of mobile food vendors with or without a permit is then $\tau(A)=\theta(A)+N_{1}(A)$.
The ratio estimator of $\theta(A)$ is
$\hat{\theta}(A)=\dfrac{N_{1}(A)\,n_{0}(A)}{n_{1}(A)}$
It is called the ratio estimator because it depends on the ratio of random variables $n_{0}(A)\,/\,n_{1}(A)$. (The total number of permitted vendors, $N_{1}(A)$, is considered fixed.) The estimator is accurate when the response rate, $p$, is approximately the same for each type, i.e.,
$p\approx\frac{n_{0}(A)}{N_{0}(A)}\approx\frac{n_{1}(A)}{N_{1}(A)}$
in which case $\hat{\theta}(A)$ is the solution for $N_{0}(A)$ in terms of $N_{1}(A)$, $n_{1}(A)$, and $n_{0}(A)$, obtained by cross-multiplication.
The main assumption of ratio estimation is that the survey data are representative in the sense that the response rate does not depend on whether a mobile food vendor has a permit. If we further assume that the response rate does not vary by subregion, we can also estimate the total number of mobile food vendors by subregion. For example, suppose the response rate in subregion $B\subset A$ is also approximately equal to $p$, e.g.,
$p\approx\frac{n_{1}(A)}{N_{1}(A)}\approx\frac{n_{0}(B)}{N_{0}(B)}$
Then we can estimate $\theta(B)$ by the subregion estimator
$\tilde{\theta}(B)=\dfrac{N_{1}(A)\,n_{0}(B)}{n_{1}(A)}$
where $\tilde{\theta}(B)$ is the solution for $N_{0}(B)$ in terms of $N_{1}(A)$, $n_{1}(A)$, and $n_{0}(B)$.
The subregion estimator is necessary when $N_{1}(B)$ is not observed and $\hat{\theta}(B)$ cannot be calculated directly. This is the case with the survey data described in Section 2.2. We observe $N_{1}(A)$, the number of permits city-wide (i.e., within region $A$), but not $N_{1}(B)$, the number of permits within subregion $B$. (We also observe $n_{0}(B)$ and $n_{1}(B)$.)
The subregion estimator $\tilde{\theta}(B)$ works by replacing the term $N_{1}(B)$ in the ratio estimator $\hat{\theta}(B)$ with the estimate
$\tilde{N}_{1}(B)=\dfrac{N_{1}(A)\,n_{1}(B)}{n_{1}(A)}$
The estimate $\tilde{N}_{1}(B)$ is also used to estimate the total number of mobile food vendors expected in subregion $B$. That is, we estimate $\tau(B)=\theta(B)+N_{1}(B)$ with
$\tilde{\tau}(B)=\tilde{\theta}(B)+\tilde{N}_{1}(B)=\dfrac{N_{1}(A)\,\left(n_{0}(B)+n_{1}(B)\right)}{n_{1}(A)}$
3.2 Model
@@
We provide a point process justification of the ratio estimation procedure described in Section 3.1. The purpose is to more closely examine the underlying assumptions and derive the standard errors of the estimates.
@@
Let $\Pi_{i}$ denote an inhomogeneous spatial Poisson process referencing the location $x\in A\subset\mathbb{R}^{2}$ of each street vendor with status $i$ at the time the survey was conducted. As in Section 3.1, we fix ideas by letting $i=0$ indicate mobile food vendors with a permit and $i=1$ indicate mobile food vendors without a permit.
@@
The Poisson process assumption may be justified by the law of rare events. New York City can be partitioned into a large number of theoretically vendable locations. Whether a vendor is located within a partition has a vanishingly small probability, such that the number of vendors in any area is well approximated by a Poisson distribution. This approximation is accurate even if these probabilities are weakly dependent, see Freedman (1974). Other justifications are possible. For example, see Section 4.2 for a justification of the Poisson process based on the theory of birth processes.
@@
so that by conditioning on $N_{1}(B)$, we arrive at the following probability model for the number of respondents with and without permits
$n_{0}(B)\sim$ $\ \text{Poisson}\big{(}p\,\theta(B)\big{)}$
$n_{1}(B)\ |\ N_{1}(B)\sim$ $\ \text{Binomial}\big{(}p,\,N_{1}(B)\big{)}$
The maximum likelihood estimates for $p$ and $\theta(A)$ are
$\hat{p}=\dfrac{n_{1}(A)}{N_{1}(A)}\qquad\text{and}\qquad\hat{\theta}(A)=\dfrac{N_{1}(A)\,n_{0}(A)}{n_{1}(A)}$
@@
Using increasing domain asymptotics, the estimator is asymptotically normal with mean $\theta(A)$ and standard error
$\text{SE}[\hat{\theta}(A)]=\theta(A)\sqrt{\dfrac{1}{p\,\theta(A)}+\dfrac{1-p}{p\,N_{1}(A)}}$
@@
Substituting $\hat{p}$ and $\hat{\theta}(A)$ for $p$ and $\theta(A)$ yields the following plug-in estimator for the standard error,
$\hat{\text{SE}}[\hat{\theta}(A)]=\dfrac{N_{1}(A)\,n_{0}(A)}{n_{1}(A)}\sqrt{\dfrac{1}{n_{0}(A)}+\dfrac{1}{n_{1}(A)}-\dfrac{1}{N_{1}(A)}}$
See Section 7.3.2 for details.
The maximum likelihood estimate for $\theta(B)$ is
$\tilde{\theta}(B)=\dfrac{N_{1}(A)\,n_{0}(B)}{n_{1}(A)}$
which is asymptotically normal with mean $\theta(B)$ and standard error
$\text{SE}[\tilde{\theta}(B)]=\theta(B)\sqrt{\frac{1}{p\,\theta(B)}+\frac{1-p}{p\,N_{1}(A)}}$
The plug-in estimator for the standard error is
$\tilde{\text{SE}}[\tilde{\theta}(B)]=\dfrac{N_{1}(A)\,n_{0}(B)}{n_{1}(A)}\sqrt{\dfrac{1}{n_{0}(B)}+\dfrac{1}{n_{1}(A)}-\dfrac{1}{N_{1}(A)}}$
To determine the subregion total $\tau(B)=\theta(B)+N_{1}(B)$, we estimate $\theta(B)$ and $N_{1}(B)$ with $\tilde{\theta}(B)$ and $\tilde{N}_{1}(B)=N_{1}(A)\,n_{1}(B)\,/\,n_{1}(A)$, yielding
$\tilde{\tau}(B)=\tilde{\theta}(B)+\tilde{N}_{1}(B)=\dfrac{N_{1}(A)\,\left(n_{0}(B)+n_{1}(B)\right)}{n_{1}(A)}=\dfrac{N_{1}(A)\,\left(n_{0}(B)+n_{1}(B)\right)}{n_{1}(B)+n_{1}(A\backslash B)}$
where $A\backslash B$ denotes the compliment of subregion $B$.
The asymptotic standard error is
$\text{SE}[\tilde{\tau}(B)]=\theta(B)\sqrt{\frac{1}{p\,\theta(B)}+\frac{1-p}{p\,N_{1}(A)}\left(1+\frac{N_{1}(B)\,N_{1}(A\backslash B)}{\theta(B)^{2}}\right)}$
and the plug-in estimator is
$\tilde{\text{SE}}[\tilde{\tau}(B)]=\dfrac{N_{1}(A)\,n_{0}(B)}{n_{1}(A)}\sqrt{\dfrac{1}{n_{0}(B)}+\left(\dfrac{1}{n_{1}(A)}-\dfrac{1}{N_{1}(A)}\right)\left(1+\dfrac{n_{1}(B)\,n_{1}(A\backslash B)}{n_{0}(B)^{2}}\right)}$
See Section 7.3.3 for details.
3.3 Data Analysis
We use the methodology described in Section 3.1 and 3.2 to estimate the number of street vendors in New York City. According to the survey data described in Section 2.2, $n_{1}(A)=349$ mobile food vendor respondents indicated they had a permit, and $n_{0}(A)=1,051$ indicated they did not. Since there are $N_{1}(A)=5,100$ permits available, it follows that the number of vendors without permits is 15,3505,100×1,051/ 349formulae-sequence153505100105134915,350$\approx$ 5,100$\times$ 1,051\,/\,34915 , 350 ≈ 5 , 100 × 1 , 051 / 349. Combined with the $5,100$ permitted vendors, the total number of mobile food vendors is approximately $20,500$.
By similar reasoning—i.e., exploiting the fact that the number of general merchandise licenses is limited for non-veterans—the number of non-veteran merchandise vendors is estimated to be approximately $1,400$. When combined with the estimated $1,000$ veteran merchandise vendors (see Section 2.1) and $20,500$ food vendors, the total number of street vendors is approximately $23,000$.
Estimates for select subregions are listed in Section 7.1, Table 1. We find that approximately one quarter of street vendors are located in the Manhattan neighborhoods of Chelsea, Clinton, and Lower Manhattan. Another quarter is located in West and North Queens. We also report the margin of error (i.e., two standard errors, half of the width of a 95% confidence interval).
We check these estimates by comparing them to a second, independent assessment conducted by the Street Vendor Project in the Bronx on May 13 and 15, 2022, in which the number of street vendors was documented through on-the-ground observations. The independent assessment identified 188 total street vendors near Fordham Road (Fordham Road BID and Street Vendor Project 2024). According to the survey data, 17 respondents indicated that they operate near Fordham Road, yielding an estimate of 152 street vendors total with a standard error of 54. We conclude this estimate is consistent with the 188 street vendors counted independently by Fordham Road BID and Street Vendor Project (2024).
4. Model Validation